Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR21 in R1 @ 15:30 pm on October 18th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Hall work

  • Shutdown activities : 28 October - 15 November
    • Mop-up of summer list and other issues?
    • SSD cryocooler system refurb (Nov 9 - Nov 15)
    • Possible test of new Linde control PC
    • Conventional magnets: Earth leakage tests at magnets (inc. synch and DSA)
    • BLOC training in synch
    • Replace Q4 Magnet-On bulb (DSA); possibly all warning lamp bulbs (synch and DSA) (Andy Gallagher, 1st Nov?)

Operations

  • MOM report and run plan
    • Straight track data tonight
    • Ramp magnets tomorrow morning
    • 6.1.4 settings from then
  • BLOC for last shift block (Oct 23 - Oct 28)
  • Magnet monitoring shifts from Oct 15 to end of running : Many (most) shifts left unfilled.
    • Need a 04:00 - 08:00 for tomorrow night

Minutes

*Present : * C Rogers, S Boyd, P Franchini, K. Long, V. Blackmore, J. Govans, P. Hanlet, L. Cremaldi, D. Sanders, M Popovic, M. Uchida

Hall work

  • DC report
  • Magnet status
    • Anomalous rampup tests were run. Nothing new was learnt. There was no reboot that preceded the ramp of M2. A reboot does not
      change the power supply parameters. The ramp of M2 is unexplained so we must maintain vigilance.
    • We still need to confirm that the power supply fixed the noise in M1.
    • Many of the communication fixes we did last week seem to have fixed the drop-out problems.
    • Run control related : Pierrick noticed that we have an absorber inside the focus coil but run control was reporting an empty one since
      that was the only thing in the CDB. This has been fixed.
    • Focus coil and spectrometers have all been resurveyed
    • FC is back online but we do need to be careful when we power it to make sure that the polarity is correct.
    • Ken : can we check that draw-wires are all connected
  • LiH absorber installation status
    • It was installed
  • Decay solenoid status
    • Craig, Vicki and Mark have been investigating the level sensor. They put a current source into the probe and it seems to
      give the right voltage, but the output signal from the readout card appears to be broken. The best bet would be to get the AMI
      controller in there instead and hook it up for readout. Trevor might be able to stop by on Friday to do this? Pierrick - this is
      a significant change and we want someone who is an expert to do this. Ken : let's see what the experts are going to do.
  • Hydrogen status
    • Insulating vacuum on hydrogen system is OK. Pressurised the absorber and there is a leak which they haven't found. The leak is not
      in the vessel but still looking.
  • Water status
    • Water is running - there is a problem with the manager which controls the two chillers. There has apparently been a spike which
      damaged the manager. Need to look at it on a shutdown job. Some tests can be done in parallel with the magnet ramp tomorrow.
  • Wednesday maintenance jobs 19th October
  • Shutdown activities : 28 October - 15 November
    • Mop-up of summer list and other issues?
    • SSD cryocooler system refurb (Nov 9 - Nov 15)
    • Possible test of new Linde control PC
    • Conventional magnets: Earth leakage tests at magnets (inc. synch and DSA)

Operations

  • MOM report and run plan
    • We stopped last week to install and things are getting going.
    • Victoria : there are a couple of things to do before we hand over for running but we are almost there.
  • Straight track data tonight
  • Ramp magnets tomorrow morning around midday
    • Data-taking to start around midday to 1 pm.
  • 6.1.4 settings from then starting with
    • Henry : can someone check the water system today. We need to check the pressure transduced. John will do that after the meeting.
  • BLOC for last shift block (Oct 23 - Oct 28)
  • Magnet monitoring shifts from Oct 15 to end of running : Many (most) shifts left unfilled.
    • Need a 04:00 - 08:00 for tomorrow night
    • We will sort this out

Updated by Nebrensky, Henry about 7 years ago ยท 9 revisions